Solve by completing the square. x^2-11x+30=0.
HELP!
Thanks!

Answer by nerdybill(7384)

x^2-11x+30=0
x^2-11x = -30
take 1/2 'b', square it and add to both sides:
x^2-11x+(11/2)^2= -30+(11/2)^2
x^2-11x+121/4= -30+121/4
(x-11/2)^2 = -120/4+121/4
(x-11/2)^2 = 1/4
take square root of both sides:
x-11/2 = +-1/2
x = +-1/2 + 11/2
x = {10/2, 12/2}
x = {5, 6}